A bottle is half full. AFter adding 275mL to the bottle it is three quarters full. How much does the bottle hold when full?
Answer by stanbon(75887)

A bottle is half full. AFter adding 275mL to the bottle it is three quarters full. How much does the bottle hold when full?
----
Let the bottle's capacity be "x":
Equation:
(1/2)x + 275 = (3/4)x
-----
(2/4)x + 275 = (3/4)x
(1/4)x = 275
---
x = 1100 mL
=================
Cheers,
Stan H.
=================